OpenStudy (anonymous):

Medal! What is the least common multiple of 70 and 30? A. 21 B. 240 C. 210 D. 10

OpenStudy (tkhunny):

Prime factorization will ALWAYS solve this sort of thing. 70 = 7*10 = 2*5*7 30 = 3*10 = 2*3*5 The least common multiple is that number whose Prime Factorization meets everyones' needs. We have from 70, 2, 5, 7. Does 30 add anything new? Yes. A 3. We are done. The Least Common Multiple is 2*3*5*7 = ??